4x + 8 = –4
Is the answer -9
Thanks for any help

Answer by longjonsilver(2297)

4x + 8 = –4
--> subtract 8 from both sides, as
4x + 8 - 8 = -4 - 8

gives
4x = -12

divide both sides by 4, as
4x/4 = -12/4

gives
x = -3

jon.